OpenStudy (anonymous):

Can someone tell me if I did this right or completely wrong!! . Verify the identity. Justify each step.  0=theta sin0 cos0 sin0 sin0 cos0 cos0 (sin0)^2 ----+---- = ----- x ---- + ----- x ------ = ------------ + cos0 sin0 cos0 sin0 sin0 cos0 sin0cos0 (cos0)^2 (sin0)^2-(cos0)^2 1 --------- = ------------------- = --------- sin0cos0 sin0cos0 sin0cos0

OpenStudy (anonymous):

tan theta + cot theta = 1/ sin theta cos theta

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

\[\large \frac{\sin \theta}{\cos \theta}+\frac{\cos \theta}{\sin \theta}=\frac{\sin^{2}\theta+\cos^{2}\theta}{\sin \theta \cos\theta}=\frac{1}{\sin \theta \cos\theta}\]

OpenStudy (anonymous):

that would be the whole thing? I dont have to break it down as much as i did?

OpenStudy (kropot72):

You could break it down as you did, it depends on the wishes of who is marking the answer.

OpenStudy (anonymous):

but my answer is still correct?

OpenStudy (kropot72):

I believe that what you have posted is correct, although it is not easy to follow unless La Tex is used.

OpenStudy (anonymous):

how do i use that? I am going to post one more question like this but now i feel like i have it expanded way to much.

OpenStudy (kropot72):

If you don't know LaTex then you can try the drawing tool here. I have to log out now, sorry.

OpenStudy (anonymous):

thats okay thank you very much